OpenStudy (anonymous):

I am not really sure if I have answered this right or not. If someone would please help I would be so thankful :) Solve 2x^2-8x=-7

OpenStudy (anonymous):

you need to use the quadratic formula! your answers should be: \[2+\frac{ 1 }{ 2 }\sqrt{2}\] \[2-\frac{ 1 }{ 2 }\sqrt{2}\]